OpenStudy (anonymous):

PLEASE HELP ill give medals & fans An oak tree that is 220 years old produced acorns at two different rates. Until it reached maturity, it produced an average of 900 acorns a year. However, once the tree reached maturity, it produced an average of 2200 acorns a year. One particular tree, Angel Oak, has produced about 289,000 acorns in its lifetime. At what age did it reach maturity? A.The tree reached maturity at 188 years. B. The tree reached maturity at 138 years. C.The tree reached maturity at 148 years. D. The tree reached maturity at 150 years.

OpenStudy (nurali):

Let the age at maturity be M 900 M + ( 220 - M) 2200 = 289000 900M + 484000 - 2200M = 289000 1300M = 195000 M = 195000 /1300 = 150

OpenStudy (anonymous):

thank you! @Nurali

OpenStudy (nurali):

My Pleasure.